OpenStudy (anonymous):

A circle has the following equation: (x+1)^2+(y-5/4)^2=49/9 Which statement is true for the circle? Its center is at the (-1, 5/4), and the radius is 7/3. Its center is at (-1, 5/4), and the radius is 49/9. Its center is at (1, -5/4), and the radius is 49/9. Its center is at (1, -5/4), and the radius is 7/3.

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

The equation of a circle is (x-h)^2 + (y-k)^2 = r^2 for some values of h,k,r constant. r is the radius and (h,k) is the location of the center of your circle. here, h = -1, k = 5/4, and r^2 = 49/9 which implies that r = 7/3 (because radii cannot be negative).

OpenStudy (anonymous):

Thank you!